OpenStudy (anonymous):

Decompose 4x2-14x + 2 devided by 4x2-1 into partial fractions (show all the steps).

OpenStudy (anonymous):

You may factor the bottom first.

OpenStudy (anonymous):

It is a difference of two squares.\[a ^{2}-b ^{2}=(a-b)(a+b)\]

OpenStudy (anonymous):

yeah i get that, im confused about the top. do i assign letters as it is or do i have to simplify?

OpenStudy (anonymous):

you will assign letters first.

OpenStudy (anonymous):

okay thanx.

OpenStudy (anonymous):

welcome. :)

OpenStudy (anonymous):

what method should i use to find A and B? after i have multiplied out.

OpenStudy (anonymous):

you mean, you already multiplied the lcd lcd on both sides? let me see your work for clarity. :))

OpenStudy (anonymous):

i got 4x2 - 14x + 2 = A(2x + 1) + B(2x - 1)

OpenStudy (anonymous):

that is correct so far.

OpenStudy (anonymous):

thanx, whtas my next step? im a lil confused by the power on the LHS i guess.

OpenStudy (anonymous):

The polynomial on the left of the equal sign must have the same coefficients as the polynomial on the right of the equal sign.

OpenStudy (anonymous):

okay,do i check that as the eqn stands oh do i have to multiply out? and if i do mutliply out will we be counting A and b as coefficients?

OpenStudy (anonymous):

i have 3 coefficients on the lhs and im wondering about the rhs?

OpenStudy (anonymous):

we can equate the coefficients of "x" first.

OpenStudy (anonymous):

the roots rights? oh i think i get it now.

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

welcome again. :)